OpenStudy (marcon):

When looking at a rational function, Bella and Edward have two different thoughts. Bella says that the function is defined at x = –1, x = 2, and x = 4. Edward says that the function is undefined at those x values. Describe a situation where Bella is correct, and describe a situation where Edward is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (marcon):

@amistre64 @NeonStrawsForever

OpenStudy (marcon):

@StudyGurl14 can you help me if you can please?

OpenStudy (studygurl14):

For the defined part, I would just do (x+1)(x-2)(x-4) = 0 For the undefined part, the easiest way I can see is to set the same thing as above up, except as a fraction, because the denominator cannot equal zero \(\large\frac{1}{(x+1)(x-2)(x-4)}=0\) Idk...something like that?

OpenStudy (marcon):

Thanks! that actually helped a lot! :D